OpenStudy (anonymous):

solve the system of simeltaneous equations : 2x+6y=15 ; 5x-8y=7 detailed answer is needed plz

OpenStudy (anonymous):

okay these two equations are said to be "solved simultaneously" that means that if you solve for x or y in one equation you can "plug it into the other equation" when you plug the x or y into the other equation you get what the y or x equals. in other words if you have two "unknows" you must have atleast two equations to solve for either variable. here the two unkowns are x and y

OpenStudy (anonymous):

to find out exactly what x and y equal you must solve these equations "simultaneously" or in other words together

OpenStudy (anonymous):

To start i have the two equations 2x+6y=15 and 5x-8y=7 you can solve either equation for either variable so randomly ill pick the first equation and ill solve that first one FOR x so 2x+6y=15 ====> x = 15/2 - 3y

OpenStudy (anonymous):

so i can rewrite what i have now 5x - 8y = 7 and x = 15/2 -3y

OpenStudy (anonymous):

from right above if you SEE that x = 15/2 - 3y AND that the top equation has an x in it. you can replace any x WITH 15/2 - 3y because x = 15/2 - 3y that is because we chose to solve for x

OpenStudy (anonymous):

so replacing the x(s) with 15/2 - 3y looks like this 5x - 8y = 7 ==> 5 (15/2 - 3y) - 8y = 7 Take a step back and look at what happend we no have an equation that has ONLY y in it. and you can solve for exactly what y is. solving for y ==> 75/2 - 15y = 7 then ==> y = 61/30

OpenStudy (anonymous):

no we have exactly what y equals that is 61/30 and we can plug this value into either of the original equations so ill plug it into the first equation 2x+6y=15 looks like 2x+6(61/30) = 15 so ==> x = 7/5

OpenStudy (anonymous):

solving equations simultaneously is this same exact thing for every simultaneous equations problem. so just repeat the major steps over and over for every different problem steps: 1 pick which equation to solve and variable to solve for and solve 2 plug into other equation 3 solve for the remaining variable 4 plug the value of the variable into either equation 5 solve for the value of the other variable

OpenStudy (anonymous):

Hey thanks very much ..

OpenStudy (anonymous):

Perhaps more straightforward, multiply one of the equations so that the coefficient of x (or y) is the same eg multiply the first equation by 5/2 to get 5x +15y = 75/2. Then subtract the new equation from the second one and the x's will disappear and give you y, then substitute y back in to get x.